### v1

# Is In-Context Learning a Type of Gradient-Based Learning? Evidence from the Inverse Frequency Effect in Structural Priming

Abstract
Large language models (LLMs) have shown the emergent capability of in-context learning (ICL). One line of research has explained ICL as functionally performing gradient descent. In this paper, we introduce a new way of diagnosing whether ICL is functionally equivalent to gradient-based learning. Our approach is based on the inverse frequency effect (IFE)—a phenomenon in which an error-driven learner is expected to show larger updates when trained on infrequent examples than frequent ones. The IFE has previously been studied in psycholinguistics because humans show this effect in the context of structural priming (the tendency for people to produce sentence structures they have encountered recently); the IFE has been used as evidence that human structural priming must involve error-driven learning mechanisms. In our experiments, we simulated structural priming within ICL and found that LLMs display the IFE, with the effect being stronger in larger models. We conclude that ICL is indeed a type of gradient-based learning, supporting the hypothesis that a gradient component is implicitly computed in the forward pass during ICL. Our results suggest that both humans and LLMs make use of gradient-based, error-driven processing mechanisms.
Is In-Context Learning a Type of Gradient-Based Learning? Evidence from the Inverse Frequency Effect in Structural Priming
Zhenghao Zhou             Robert Frank             R. Thomas McCoy Yale University 370 Temple Street New Haven, CT 06511 {herbert.zhou, robert.frank, tom.mccoy}@yale.edu
1 Introduction
To what extent do humans and language models use similar processing mechanisms? This question is of interest to both Artificial Intelligence researchers and cognitive scientists. Language models and human learners have some substantial differences: human learners often display a flexible learning ability to adapt to new examples, while language models typically require massive amounts of training data and a large number of parameters to exhibit human-like performance. Yet recent pre-trained large language models (LLMs) have shown the capacity for in-context learning (ICL): they adapt to specific tasks with a few demonstration-answer pairs provided as prompts in the context window without any parameter updates (Brown et al., 2020). This intriguing, emergent capability could provide a way to bridge the divide between language models and human learners: perhaps ICL is a processing mechanism that, like humans, can flexibly adapt to new demands.
Among the body of research into the sources and interpretations of the ICL capability, one line of work has aimed to deepen the theoretical understanding of ICL by offering functional interpretations of ICL as a kind of implicit gradient descent during inference. Garg et al. (2022), Zhang et al. (2024), and Ahn et al. (2024) have shown that standard Transformers (Vaswani et al., 2017) can be trained to implement learning algorithms for linear regressions under ICL-based training objectives. Von Oswald et al. (2023) have demonstrated that Transformer models, with appropriate choices of parameters, can process in-context demonstrations in a way that is functionally equivalent to performing gradient updates on the same demonstration examples. Dai et al. (2023) provided a mathematical construction showing the dual form between Transformer attention and gradient descent and interpreted ICL as a meta-optimization process that performs implicit fine-tuning. However, Shen et al. (2023) observe that the importance of these theoretical demonstrations is limited in that they do not take ICL to be an emergent property, instead assuming a training objective that optimizes for ICL. They therefore deviate from actual LLMs pre-trained with natural data with a language modeling training objective. Indeed, Shen et al. found inconsistencies between ICL and gradient descent in real models, and left the equivalence between ICL and gradient descent an open question.
In this paper, we aim to better characterize what kind of learning mechanism ICL is by drawing a connection between ICL and human learning mechanisms. Specifically, we examine the hypothesis that ICL functionally performs gradient-based fine-tuning (e.g., gradient descent) by empirically investigating a weaker claim with off-the-shelf LLMs evaluated with natural language data: whether ICL is a type of gradient-based—i.e., error-driven—learning such that a gradient component is implicitly computed during the forward pass. We approach this question by treating ICL as a processing mechanism of LLMs and borrowing insights from methods of studying processing mechanisms in humans: we examine to what extent LLMs show the inverse frequency effect (IFE), a phenomenon in the human structural priming paradigm (Branigan and Pickering, 2017) that has been argued to require a gradient-based processing mechanism in humans, namely implicit learning (e.g., Chang et al., 2006). We study the linguistic phenomenon of the dative alternation and demonstrate that LLMs show a robust IFE under standard fine-tuning (which involves explicit gradient-based learning), and varying degrees of the IFE under the ICL setting, with larger models showing a stronger IFE. We conclude that, at least in the cases we studied, ICL is indeed a gradient-based learning mechanism.
Our study has implications for both NLP/machine learning (1 and 1) and linguistically-motivated analysis of LLMs (1 and 1):
[…]
. We find evidence that ICL in off-the-shelf LLMs can be viewed as a form of gradient-based learning.
[…]
. We generalize the notion of ICL beyond the standardly assumed prompt format of input-output pairs, establishing a connection between priming and prompting.
[…]
. We show that LLMs qualitatively display an important property of human language processing, namely the IFE in structural priming.
[…]
. While most human-LLM comparisons focus on representations, our experiments go one step further by analyzing the processing mechanisms used by LLMs.
Overall, our results suggest that error-driven learning is an aspect of processing that is shared between humans and LLMs.
2 Background and Related Work
In this section, we lay out the building blocks necessary for motivating why we use the IFE to diagnose the gradient-based nature of ICL. Our experimental approach is formally described in Section 3.1.
Figure 1: Reasoning behind our current study. Abbreviations: LLMs = large language models; ICL = in-context learning; IFE = the inverse frequency effect (the tendency for error-driven learners to show larger updates when trained on a low-frequency example than a high-frequency one; see text).
2.1 Structural Priming in Psycholinguistics
Structural priming refers to the phenomenon that speakers tend to reuse recently encountered syntactic structures Bock (1986). For example, after encountering a double object (DO) sentence (e.g., Alice gave Bob a book), speakers tend to produce another DO structure (e.g., The student sent the professor a letter) rather than a prepositional dative (PD) structure (e.g., The student sent a letter to the professor). Structural priming has been interpreted as an adaptation mechanism, where speakers adapt lexical and syntactic predictions to the current context Jaeger and Snider (2013), similar to the way in which LLMs adapt their outputs on the basis of prompts. Examples of priming typically involve two sentences; we will refer to them as the prime sentence and the target sentence, where the prime sentence comes first and influences predictions about the target sentence.
One important aspect of structural priming is the inverse frequency effect Jaeger and Snider (2008); Bernolet and Hartsuiker (2010); Kaschak et al. (2011): less preferred syntactic alternatives (measured by the relative frequency in the speaker’s experience against their counterparts) cause stronger overall priming than more preferred structures. This can be seen by looking at sentences involving verbs with gradient structural preferences, or verb biases (or alternation biases; see Hawkins et al., 2020 for a systematic investigation of verb biases in neural models). The verb give allows both prepositional dative (PD) (The doctor gave a book to the judge) and double object (DO) (The doctor gave the judge a book) structures, but is biased toward (i.e., occurs more often with) DO in English. Under the IFE, a “prime” sentence involving give in its (less preferred) PD structure will cause a greater priming effect than a DO sentence, i.e., a PD prime will increase the probability of a PD occurring more than a DO prime will increase the probability of a DO occuring. The strength of PD priming (i.e., the increase in the probability of a PD target given a PD prime) inversely correlates with the expectation of a PD prime, determined by its verb bias Bernolet and Hartsuiker (2010).
Two mainstream theories have been proposed to account for structural priming. Transient activation theory Pickering and Branigan (1998) claims that the activation of structural representations from the prime persists for a short time (in working memory), such that the structural information has a higher probability of being reactivated on the next relevant opportunity. The current form of transient activation theory does not account for the IFE because the amount of residual activation is independent from verb biases and does not involve any error-driven mechanism. In contrast, implicit learning theory Chang et al. (2006) claims that humans implicitly learn and dynamically adjust probabilistic information about different structures (including verb biases) from experience (in long-term memory) and use such information to predict the form of prime sentences. Crucially, under standard theories of learning, the update performed by the learner is error-driven, such that a larger update is performed in situations where the learner’s predictions are farther from the truth. In the context of priming, this would mean that priming strength is determined by the difference between the learner’s predictions and the actual prime sentence: the less the learner expects the observed prime structure, the larger the gradient is, resulting in a larger priming strength. Therefore, the implicit learning theory—unlike transient activation—predicts the IFE. The two theories are not mutually exclusive and can co-exist to account for priming, a framing known as the dual mechanism account Tooley and Traxler (2010).
In this study, we assume the correctness of the arguments from psycholinguistics that state that only error-driven learning mechanisms could predict the IFE. Therefore, by examining whether LLMs show the IFE in the ICL setting, we can infer whether some type of gradient component is computed in the forward pass without explicit weight updates, which informs us about whether there is an implicit gradient-based component in ICL.
2.2 Connections among Distributional Properties of Pre-Training Data, Priming, and In-Context Learning
Alongside the theoretical accounts of the basis of ICL mentioned earlier, another line of research explains the origin of ICL through the distributional properties of the pre-training data. Chan et al. (2022) showed that ICL emerges when the training data exhibits particular distributional properties (such as burstiness, in which items appear in clusters rather than being uniformly distributed over time). Hahn and Goyal (2023) argued that ICL emerges from the compositional structures found in the pre-training data under the standard next-token prediction objective. Chen et al. (2024) found that parallel structures in the pre-training data give rise to ICL capability in LLMs. They defined parallel structures as pairs of phrases following similar templates in the same context window and found that removing parallel structures in the pre-training data significantly reduces LLMs’ ICL accuracy. Chen et al. also pointed out that despite the fact that the pre-training data is not formatted strictly as in-context prompts, i.e., input-output pairs, the naturalistic data often contains phrases following similar templates. Those phrase pairs could be conceptualized as in-context examples of implicitly defined, less structured shared “tasks”, such as copying n-grams, repeating syntactic constructions, and producing related types of world knowledge.
As structural priming is a well-attested phenomenon in humans, it is reasonable for us to hypothesize that structural priming is a factor that shapes the distribution of LLM pre-training data since humans tend to produce abundant parallel structures in naturalistic settings. For this reason, we view the repeated structures from structural priming as a case of the parallel structures in Chen et al.’s (2024) sense. Inspired by this data-centric perspective, we do not take ICL to necessarily involve explicit demonstration-answer pairs for specific tasks, as is typically understood in the literature. Instead, we conceptualize ICL as a more generalized notion that involves a sensitivity to parallelism through generic next-token prediction: any text in the context window will influence the model’s logits toward preferring similar text. This generalized notion of ICL, namely, the contextual influence of prompts in the context window, is analogous to priming in humans, as we elaborate in Section 3.1.
2.3 Structural Priming in Neural Language Models
Structural priming has been proposed as a means of probing the abstract mental representations of structural information in humans that define the notion of similarity underlying priming (Branigan and Pickering, 2017). Previous works have adopted this paradigm for probing learned linguistic representations in neural networks. van Schijndel and Linzen (2018) and Prasad et al. (2019) show that LSTMs (Gulordava et al., 2018) are capable of adapting to syntactic structures under fine-tuning. Specifically, these researchers fine-tune model weights on prime sentences and test target sentence probabilities in the updated model. This is a direct analog of the implicit learning account of structural priming as it involves weight updates. Recently, Sinclair et al. (2022) have shown that the GPT2 family Radford et al. (2019) shows robust structural priming without the use of explicit weight updates; when a prime sentence is concatenated to its corresponding target sentence, we see an increase in the probability the model assigns to the target sentence as compared to without a prime. [^1] Sinclair et al. (2022) have also demonstrated that the GPT2 models showed the lexical boost effect, another well-attested sub-phenomenon of structural priming, which is not our main focus here. Other works have demonstrated crosslingual structural priming in large language models Michaelov et al. (2023), suggesting that structural priming is robustly detected in LLMs. Jumelet et al. (2024) study in considerable details the linguistic factors the give rise to structural priming in current state-of-the-art LLMs.
In sum, there is a large body of work demonstrating that LLMs, like humans, show structural priming. This sets the stage for our current study of investigating the processing mechanisms underlying priming.
3 Current Study
Figure 2: An overview of our experimental design.
3.1 Overview of Our Approach
We first clarify our conceptualization of ICL. As stated in Section 2.2, instead of adopting the notion that ICL requires demonstration-answer pairs of a task as prompts in the context window, here we propose that any text in the context window will predispose the model toward producing similar text in the future. This text could take the form of demonstration-answer pairs, as is the case in the best-known examples of ICL, but it does not need to: it could be any type of text whose structure is then preferred by the model in subsequent predictions. Under this generalized framing, ICL can produce structural priming: a model that observes a sentence with a particular syntactic structure (say, the DO structure) can use ICL to replicate that syntactic structure, even though the sentence cannot be cleanly viewed as a demonstration-answer pair. That is, the less structured, implicitly defined “task” encoded by the DO sentence used as the prompt could be interpreted as “producing another sentence following the DO structural template exemplified in the prompt.”
Our research question is whether a gradient component is implicitly computed during the forward pass of processing the prompt in the generalized ICL setting. We investigate this question by testing whether LLMs show the IFE in the ICL setting. [^2] In work that was done concurrently with the present paper, Jumelet et al. (2024) provide evidence that LLMs show the IFE, corroborating this finding of our paper. The two papers differ in several ways. First is the main motivation: we use the IFE as a diagnostic for characterizing the type of mechanism that underlies ICL, while Jumelet et al. study the IFE for purposes of understanding how similar LLM structural priming is to human structural priming. A second difference stems from the analyses that are run: the current paper defines an explicit metric for the IFE in Section 3.6 based on each prime verb’s individual verb bias, while Jumelet et al. use a linear mixed model that involves a different metric of the verb biases (i.e., structural preference) to investigate the predictive power of various factors on the LLMs’ priming behaviors. Finally, the two papers differ in the form of evidence that they use for detecting the IFE: we show direct evidence for the IFE in LLMs through experimental design that isolates the effect of frequency, while Jumelet et al. present indirect evidence for the IFE in LLMs by analyzing which predictors in a linear mixed model have a significant effect: they show that the sentence-level surprisal and the structural preference are predictive of the priming effects, and they treat these two factors as evidence for the IFE in LLMs. As is illustrated in Figure 1, given that (i) it has been argued by psycholinguists that only error-driven learning mechanisms can give rise to the IFE; (ii) processing the prime sentence in the context window conditions the probability of the target sentence in the generalized notion of ICL; and (iii) standard structural priming in the ICL setting has been robustly observed , we hypothesize that the strength of the IFE will positively correlate with the strength of the ICL capability of LLMs: the stronger the ICL capability is, the better the gradient will be computed in the forward pass, leading to a stronger IFE.
Specifically, we simulate structural priming across LLMs of various sizes with the two modes mentioned in Section 2.3. As is illustrated in Figure 2, the Fine-Tuning mode fine-tunes the LLM’s parameters on a single prime sentence, and the updated model is used to infer the probability of the target sentence. The Concatenation mode instead uses (our generalized notion of) ICL, where the prime sentence is directly concatenated with the target sentence, using the prime sentence as the prompt in the context window, and the probability of the target sentence is measured. The Fine-Tuning mode serves as a sanity check that LLMs are able to show the IFE when there is explicit error-driven, gradient-based learning. It sets the stage for our main focus: using the Concatenation mode to diagnose whether ICL involves implicit gradient-based learning.
3.2 Corpus
We adapted the Core Dative Prime-LM Corpus from Sinclair et al. (2022) to create our dataset. We briefly introduce the relevant properties of the corpus and refer the readers to the original paper for details. The dative corpus consists of sentences in two forms:
[…]
. DO: DPsubj V DPiobj DPdobj
e.g., A girl bought a guy a coffee.
[…]
. PD: DPsubj V DPdobj Prep DPiobj
e.g., A girl bought a coffee for a guy.
Each DP is a determiner with a common noun (120 distinct nouns in total). The corpus was constructed in a way that controlled for the degree of semantic association and lexical overlap between prime and target sentences, and sentences are semantically plausible as the ditransitive verbs were manually labeled with their verb frames.
Since our goal is to study the IFE, we will need to ensure variation in the occurrence of verbs with different verb biases in the the prime and target sentences. To do this, for each prime sentence involving one of 22 possible verbs, we sampled 50 target sentences with no lexical overlap. This means that the verbs in prime and target sentences are always distinct. Each prime-target pair can be instantiated structurally in 4 different ways, depending on whether the target and prime sentences are PD or DO: $t_{\text{PD}}|p_{\text{PD}}$, $t_{\text{PD}}|p_{\text{DO}}$, $t_{\text{DO}}|p_{\text{PD}}$, $t_{\text{DO}}|p_{\text{DO}}$ (i.e., target sentence $t$ conditioned on prime $p$). This resulted in 92400 prime-target pairs. [^3] In this paper, we use $t$ and $p$ for individual target and prime sentences, $T$ and $P$ for sets of target and prime sentences, and $\mathcal{P}$ for probability. An example of $t_{\text{PD}}|p_{\text{DO}}$ is “A doctor brought a chief a plate. The secretary drew the card for the band.”
We also created an alternative dataset of the same size by replacing all indirect object DPs with a pronoun. [^4] Details of the set of pronouns and their relative probabilities are in Appendix A. This was motivated by a corpus parse [^5] In order to estimate the verb biases represented in the training corpus of GPT2 models, we parsed a fragment (around 160 million tokens) of the OpenWebText corpus (Gokaslan and Cohen, 2019) with the Python package spaCy (Honnibal et al., 2020) to get an estimate of the DO vs. PD ratio for each verb. We found that the verb biases from the corpus are less well-represented in GPT2 models, motivating our decision to estimate verb biases by using model judgments rather than corpus statistics; see Section 3.4. that showed that the most common indirect objects in DO sentences are animate pronouns, suggesting that it may be important to consider pronouns in analyses of DO vs. PD sentences. The importance of animacy has also previously been noted by Bresnan et al. (2007). The presence and absence of pronouns lead to different verb biases for LLMs, which affect their IFE behaviors. We will return to this point in the discussion.
3.3 Language Models
We considered a set of Transformer models that have been claimed to show ICL capabilities to various extents (Lee et al., 2024):
  • GPT2 Radford et al. (2019) in three of its sizes (small, medium, large), with 85M, 302M, and 708M parameters, respectively. All versions were loaded from the package transformerLens Nanda and Bloom (2022).
  • LLAMA2 Touvron et al. (2023) in three versions: 7b (6.5B parameters), 7b-chat (6.5B parameters), and 13b (13B parameters). All versions were loaded using the Huggingface transformers library Wolf et al. (2020).
  • GPT3-base Brown et al. (2020) with the davinci-002 version (175B parameters), accessed via the OpenAI API.
The models vary in size, and correspondingly, in their ICL capabilities. We predict that there will be a stronger IFE as size increases. [^6] We also tested the LSTM models from (Gulordava et al., 2018) with the Concatenation mode and found no evidence of structural priming.
Figure 3: A demonstration of the IFE: a stronger priming effect of a DO prime is predicted as PD-bias increases. Note that, because the prime and target sentence differ in structure, the target sentence’s log probability is lower (more negative) when the effect of priming is larger. The numerical values of the primed log probabilities are hypothetical, for illustration purposes only.
3.4 Quantifying Verb Biases
The verb bias for a specific verb is the likelihood of producing structure $X$ compared to the alternative structure $Y$. In human experiments, baseline verb biases are estimated as the ratio of the number of times one structure occurs over the sum of both structures; such occurrences can be counted either in natural production experiments or in corpus searches Zhou and Frank (2023). Here, we compute a verb bias for each verb analogously as the ratio of the model probability of one structure over the sum of the probabilities of both structures. The probability of a sentence $s$ is the product of probabilities assigned by LLMs to each token $w_{i}$: $\mathcal{P}(s)=\prod_{i}\mathcal{P}(w_{i}|\overline{w}_{<i})$. [^7] In practice, we took the sum of the log probabilities assigned by LLMs to each token in the target sentence, which is equivalent to taking the log of the product of raw probabilities. This measures how likely it is for the model to see or produce this sentence. Then, given a set of sentences $\mathcal{S}_{V}$ with ditransitive verb $V$, where each sentence $t_{\text{PD}}$ with structure PD always has its counterpart $t_{\text{DO}}$ in the opposite structure of DO (see 3.2 and 3.2), the $\mathbf{PD}$-bias of verb $\mathbf{V}$ is the mean normalized probability of sentences in structure $PD$:
$bias(V,PD)=\frac{1}{|\mathcal{S}_{V}|}\sum_{t_{\text{PD}}\in\mathcal{S}_{V}}\frac{\mathcal{P}(t_{\text{PD}})}{\mathcal{P}(t_{\text{PD}})+\mathcal{P}(t_{\text{DO}})}$ (1)
The DO-bias of verb $V,$ i.e., $bias(V,DO)$, is defined analogously. We compute verb bias separately for each model, and use a model’s biases in our assessment of whether it shows the IFE.
3.5 Simulating Structural Priming
As stated in Section 3.1, we use two modes to simulate structural priming. Following van Schijndel and Linzen (2018), for the Fine-Tuning mode, we update the parameters by fine-tuning the model on a single prime sentence with learning rate $1e^{-5}$ for 10 epochs (see the full fine-tuning details in Appendix B), and we use the updated model to do inference on the target sentence. Following Sinclair et al. (2022), for the Concatenation mode, we condition a target sentence on a prime sentence through directly concatenating them, separated by a period, without any weight updates.
The probability of the target sentence after priming is the product of probabilities assigned to its tokens: $\mathcal{P}(t_{X}|p_{X})=\prod_{i}\mathcal{P}(t_{X_{i}}|p_{X},t_{X_{<i}})$. Under the standard priming effect, the probability of the target sentence $t_{X}$ should be greater after being primed by a ditransitive prime sentence of either structure: $\mathcal{P}(t_{X}|t_{X})>\mathcal{P}(t_{X})$ and $\mathcal{P}(t_{X}|p_{Y})>\mathcal{P}(t_{X})$. Being primed by the same structure induces a larger probability increase than being primed by the opposite structure: $\mathcal{P}(t_{X}|p_{X})>\mathcal{P}(t_{X}|p_{Y})$.
3.6 Predictions about the Inverse Frequency Effect
Recall that the IFE states that the priming strength of structure $X$ inversely correlates with the prime verb’s $X$-bias. That is, the degree to which the target production’s probability deviates from the baseline is determined by the prime verb. Consider, for example, the case in which the prime sentence has a DO structure and the target sentence has a PD structure (other combinations of structures work analogously). For each prime verb $V$, we computed the PrimeBias for the PD target structure given a DO prime sentence, where PrimeBias is defined as the normalized target probability primed by this verb over a set of target sentences:
$\begin{split}PrimeBias(\text{PD}|\text{DO},V)&=\frac{1}{|T_{\text{PD}}|\cdot|P_{\text{DO}}^{V}|}\sum_{t_{\text{PD}}\in T_{\text{PD}}}\sum_{p_{\text{DO}}^{V}\in P_{\text{DO}}^{V}}\\ &\frac{\mathcal{P}(t_{\text{PD}}|p_{\text{DO}}^{V})}{\mathcal{P}(t_{\text{DO}}|p_{\text{DO}}^{V})+\mathcal{P}(t_{\text{PD}}|p_{\text{DO}}^{V})}\end{split}$
As shown in Figure 3, the IFE predicts that with a PD target and DO prime sentence, as the prime verb $V$’s PD-bias increases, the prime sentence is less expected, resulting in a larger priming strength towards the DO direction in the target production, i.e., a smaller $PrimeBias(\text{PD}|\text{DO},V)$ value. Similarly, as PD-bias increases, a PD prime sentence will result in a smaller priming strength towards the PD direction in target production, i.e., again a smaller $PrimeBias(\text{PD}|\text{PD},V)$ value. Therefore, when plotting $PrimeBias(\text{PD}|\text{DO},V)$ and $PrimeBias(\text{PD}|\text{PD},V)$ against increasing verb biases and fitting a line with linear regression, the IFE predicts negative slopes for both plots. Moreover, standard priming predicts that $PrimeBias(\text{PD}|\text{PD},V)$ should have a higher intercept than $PrimeBias(\text{PD}|\text{DO},V)$ since the former increases the probability of $T_{\text{PD}}$ more than the latter. [^8] The other two conditions, namely $T_{\text{DO}}|P_{\text{PD}}$ and $T_{\text{DO}}|P_{\text{DO}}$, are mathematically guaranteed to have slopes that are exactly opposite from $T_{\text{PD}}|P_{\text{PD}}$ and $T_{\text{PD}}|P_{\text{DO}}$, respectively, and the intercept for each should be 1 minus the intercept of its counterpart. We do not plot these other two conditions because these mathematical relationships mean that they provide no information that is not already provided by the plots we do show.
4 Results and Analysis
Figure 4: GPT2-small shows a robust IFE under the Fine-tuning mode; the negative slope in these plots is the signature of the IFE. Both the $T_{\text{PD}}|P_{\text{PD}}$ (left) and $T_{\text{PD}}|P_{\text{DO}}$ (right) conditions have negative slopes, and the $T_{\text{PD}}|P_{\text{PD}}$ has a higher intercept than $T_{\text{PD}}|P_{\text{DO}}$.
For each model and for each prime verb, we plotted $PrimeBias(\text{PD}|\text{PD},V)$ and $PrimeBias(\text{PD}|\text{DO},V)$ against increasing verb biases and used linear regression to find the relationship between priming strength and verb biases. We report the R-squared ($R^{2}$) coefficient and the root mean squared error (RMSE) to assess the significance of the fitted lines.
Figure 5: The IFE across models of different sizes in the WithPronoun condition under the Concatenation mode.
Models With Pronoun PDPD_slope PDPD_intercept PDPD_$\mathbf{R^{2}}$ PDPD_RMSE DOPD_slope DOPD_intercept DOPD_$\mathbf{R^{2}}$ DOPD_RMSE GPT2-small True 0.011 0.370 0.014 0.020 -0.007 0.278 0.008 0.017 GPT2-small False 0.014 0.746 0.024 0.016 0.006 0.653 0.003 0.019 GPT2-medium True -0.013 0.351 0.015 0.023 -0.026 0.256 0.107 0.016 GPT2-medium False -0.023 0.748 0.067 0.017 -0.035 0.590 0.060 0.027 GPT2-large True 0.011 0.330 0.017 0.019 -0.037 0.241 0.173 0.018 GPT2-large False -0.003 0.698 0.001 0.018 -0.020 0.487 0.026 0.024 LLAMA2-7b True -0.020 0.392 0.073 0.015 -0.086 0.229 0.645 0.013 LLAMA2-7b False -0.026 0.807 0.046 0.019 -0.111 0.627 0.149 0.042 LLAMA2-7b-chat True -0.012 0.413 0.019 0.018 -0.095 0.263 0.587 0.017 LLAMA2-7b-chat False -0.013 0.788 0.007 0.024 -0.102 0.605 0.107 0.044 LLAMA2-13b True -0.059 0.434 0.323 0.018 -0.099 0.256 0.760 0.011 LLAMA2-13b False -0.066 0.859 0.160 0.019 -0.177 0.685 0.224 0.042 davinci-002 True -0.078 0.403 0.570 0.013 -0.078 0.223 0.662 0.011 davinci-002 False -0.064 0.851 0.172 0.020 -0.145 0.632 0.257 0.035
Table 1: The slope, intercept, $R^{2}$, and RMSE of the fitted lines for each condition under the Concatenation mode. Conditions with both negative slopes are bold (which suggests capturing the IFE), and $R^{2}$ scores higher than $0.5$ are bold (which means a more significant fitted line). PDPD means that the prime and target sentence both have a PD structure, while DOPD means that the prime sentence has a DO structure and the target sentence has a PD structure.
4.1 Fine-tuning Mode
We applied the Fine-tuning mode to GPT2-small. [^9] We did not carry out this mode for larger models because of the substantial computational resources they require: each priming instance requires a separate fine-tuning process. However, this is unproblematic for our conclusions, given that the Fine-tuning mode is expected to show the IFE in all cases, given its explicit gradient updates. As is shown in Figure 4, the $T_{\text{PD}}|P_{\text{PD}}$ condition has a larger intercept than the $T_{\text{PD}}|P_{\text{DO}}$ condition, suggesting that the Fine-tuning mode is able to capture standard structural priming. Further, we observe negative slopes in both plots, suggesting that the Fine-tuning mode is able to capture the IFE. The $T_{\text{PD}}|P_{\text{DO}}$ condition has a higher $R^{2}$ score of $0.75$, demonstrating a stronger IFE than the $T_{\text{PD}}|P_{\text{DO}}$ condition.
Overall, these results show that even the smallest model we analyzed shows the IFE under explicit gradient-based weight updates, which passes the sanity check by confirming that LLMs are capable of showing the IFE with explicit gradient-based weight updates. We now turn to the crucial case of ICL, to see whether the IFE arises in LLMs even without explicit weight updates.
4.2 Concatenation Mode
We applied the Concatenation mode to all models. We only show one plot for each of the three types of models (i.e., showing just one variant of that model type and showing only the WithPronoun condition) and report the full results (with and without pronouns across all model variants) in Table 1. As shown in Figure 5, for all models across all conditions, the $T_{\text{PD}}|P_{\text{PD}}$ intercept is greater than the $T_{\text{PD}}|P_{\text{DO}}$ intercept, showing the standard structural priming effect, which is consistent with our prediction. The RMSE score is less than $0.04$ for all conditions, suggesting a significant predictability of the fitted lines to the data points. For the IFE, we found that all three sizes of GPT2 failed to show the IFE, as the slopes are either positive or close to zero. This suggests that, in GPT2, priming strength is not correlated with verb bias under our metric. All three LLAMA2 models showed negative slopes in both cases, which is consistent with the IFE. However, only in the WithPronoun $T_{\text{PD}}|P_{\text{DO}}$ condition are the $R^{2}$ coefficients constantly greater than $0.5$ across the three models, [^10] Given no consensus on standard $R^{2}$ score thresholds, we picked this criterion by default. suggesting that LLAMA2 displays the IFE but in a noisy way that is not fully robust. Finally, for GPT3, both $T_{\text{PD}}|P_{\text{PD}}$ and $T_{\text{PD}}|P_{\text{DO}}$ under the WithPronoun condition have an $R^{2}$ coefficient greater than $0.5$, while $R^{2}$ is not greater than 0.5 in either case in the NoPronoun condition.
Therefore, besides corroborating the finding from prior work that LLMs show structural priming effects, the current results suggest that, in general, larger models tend to show a stronger IFE than smaller models, meaning that IFE strength correlates with ICL capability (since larger models also typically perform better at ICL than smaller models). Assuming that ICL capability correlates with LLM size, given the currently observed pattern, we predict that larger models such as GPT4 should show an even stronger and more significant IFE.
4.3 The Distinction between the WithPronoun vs. NoPronoun Conditions
As shown in Table 1, the majority of cases with an $R^{2}$ score above $0.5$ are the WithPronoun $T_{\text{PD}}|P_{\text{DO}}$ cases. Why the IFE arises more strongly in the WithPronoun condition than the NoPronoun condition remains unclear. One major difference between these conditions lies in the default verb biases: as shown in Figure 6, GPT3 shows an overwhelming bias towards PD without pronouns but a reverse pattern favoring DO with pronouns. This pattern holds across all models and is consistent with statistics obtained from our corpus parse. One possible conclusion is that the most common indirect object DPs in the DO sentences are animate pronouns, causing the model to assign a higher probability to pronoun sentences. However, it still remains puzzling why and how these differences in verb biases could lead to the IFE having different levels of significance across the two conditions.
Figure 6: Comparison of PD biases with (left) and without (right) pronouns for GPT3. As is shown in Equation 1, a high PD-bias means a larger proportion of probability mass is assigned to the PD structure than the DO structure in LLMs.
5 Discussion and Conclusion
Evidence that ICL is implicitly a gradient-based learning mechanism
We started with the question of whether ICL could be a processing mechanism of LLMs that can capture the flexibility and adaptability of human learning mechanisms. One type of learning mechanism that supports flexible adaptation to data is gradient descent. Motivated by previous theoretical proposals that ICL could involve functional gradient descent or implicit fine-tuning, we attempted to better characterize what kind of learning ICL is in actual models. We focused on one particular aspect of ICL: whether it involves an implicit gradient component during the forward computation. We differ from previous approaches for analyzing ICL in that we test standard LLMs with natural language data. Having established a connection between ICL and human structural priming, we used the IFE to diagnose whether LLMs behave as if they are implicitly performing gradient descent when processing a prime sentence. We found that LLMs do indeed display the IFE in many cases, particularly in larger models. These findings support the hypothesis that a gradient computation is implicitly performed in the forward pass of ICL. Differences between models suggests that ICL as a gradient-descent-like process arises only in larger models. More speculatively, these results raise the possibility that gradient-based learning might be a crucial property that enables generalization from a small number of samples, an ability that is shared by human learners and (to some extent) LLMs performing ICL. Our study not only provides behavioral results that align LLMs’ behaviors with human behaviors on structural priming at the processing mechanism level, but also demonstrates the possibility of studying the nature of ICL with off-the-shelf pre-trained LLMs and with naturalistic data.
ICL as a consequence of language modeling?
ICL is typically understood as involving demonstration-answer pairs in the prompt. Inspired by data-centric views that explain ICL from the distributional properties of pre-training data, we proposed a generalized notion of ICL that is sensitive to general parallelisms. As a result, any text in the prompt could serve as an implicitly defined “task” of following the template provided in the context and generating a parallel structure. Therefore, ICL could be viewed as a side product of the general propensity for structural parallelism in the language modeling task. We leave this perspective for future investigation.
Future Directions
In this study, we focused on one single “task”, namely selecting DO versus PD sentence structures. If our reasoning is correct, such that it is indeed the implicit gradient component of the ICL that results in LLMs’ capability of capturing the IFE, then we predict that the IFE diagnostics could be generalized to other ICL tasks, even non-linguistic tasks. Future work could extend our current method to additional ICL tasks. Finding the IFE on a wider range of tasks would strengthen the claim that ICL is driven by implicit gradient computation, while not observing the IFE on other tasks would indicate that ICL only sometimes displays a gradient-descent-like character.
Limitations
Behavioral versus Mechanistic Accounts
Although ICL is generally identified as a phenomenon at the behavioral level, having an explanation at the mechanistic level is desirable since it brings greater interpretability and contributes to more concrete theory building. Our current study, despite using real pre-trained models and naturalistic data, remains at the behavioral level and is empirical in nature. Given our current contribution of establishing a connection between ICL and human priming and using the IFE as a way to diagnose whether ICL works in a gradient-descent-like manner or not, future work could improve our understanding by incorporating techniques from mechanistic interpretability to explain our current finding at the mechanistic level.
Examining the IFE in Other Models
As ICL capability is often argued to scale with model size, we predict in Section 4.2 that the IFE effect will be more robust in larger models. Although the difference in the IFE behavior between GPT2 and GPT3-base is substantial, we do not believe that we have observed a saturation of the IFE because the $R^{2}$ values that have observed are far from the ceiling level of 1.0. GPT3-base is currently the biggest model on which we have access to the logit predictions, but we believe the same behavioral test could be applied to larger models in order to verify our prediction.
Acknowledgments
We would like to thank the anonymous reviewers from the ESSLLI Student Session 2024 and AMLaP 2024 for their thoughtful comments, which helped us refine this paper. We would also like to thank the members of the Yale Linguistics Department for helpful discussion. Any errors remain our own.
Appendix A Finding pronoun probabilities in the OpenWebText corpus with spaCy
As is mentioned in Section 3.2, we constructed the WithPronoun version of the corpus in order to investigate the impact of animacy of the indirect object on the verb biases. To do this, we approximated the distribution of the natural occurrence frequencies over the set of English pronouns in dative alternation sentences from a fragment of the OpenWebText corpus (Gokaslan and Cohen, 2019), which is a reasonable proxy for the closed WebText corpus that was used to train the GPT2 models.
We parsed a fragment (around 160 million tokens) of the corpus with spaCy (Honnibal et al., 2020). Specifically, we used the en_core_web_trf specification of the spaCy model, and we identified the set of dative alternation sentences by doing dependency parsing on each sentence. Then, we counted the frequencies of the set of English pronouns that occurred as the indirect object of the ditransitive verb. The list of pronouns and their frequencies are presented in Table 2, sorted by frequency.
Pronoun Frequency
you 4621
me 2962
us 2959
him 2210
them 1847
it 1297
her 738
Table 2: The respective frequencies of the English pronouns occurring as the indirect object of ditransitive sentences in a fragment of the OpenWebText corpus.
To convert the existing dative alternation priming corpus to the WithPronoun version, we replaced the indirect object of every sentence in the existing corpus with one of the pronouns through random sampling according to their respective relative frequencies.
Appendix B Fine-tuning details
As is presented in Section 3.5, to simulate structural priming in the Fine-tuning mode, we fine-tuned a pre-trained GPT2-small model on every prime sentence and used the updated model to do inference on the target sentences.
We loaded the pre-trained GPT2-small model from the TransformerLens (Nanda and Bloom, 2022) package and used the train function from TransformerLens to do fine-tuning. To avoid catastrophic forgetting during fine-tuning, we applied a regularization term to the loss function for gradient descent. We randomly sampled a fixed set of 5000 adjacent tokens from the OpenWebText (so that it resembles the distribution of the pre-training data) and computed the loss on them of the pre-trained GPT2-small model. Then, at each step during fine-tuning, we added to the loss term the squared difference between the current loss and the raw (pre-trained) loss of the model on these 5000 tokens, scaled by a coefficient $\lambda=0.8$. We found that this regularization term helped keep the model stable during fine-tuning on a single sentence.
Parameter Value
number of epochs $10$
batch size $1$
learning rate $1e^{-5}$
optimizer AdamW
lambda 0.8
Table 3: Hyperparameters used as the training configuration for the Fine-tuning mode of structural priming on GPT2-small.
We did a hyperparameter search and chose the set of parameters in Table 3. We used the default values from TransformerLens for the rest of the relevant hyperparameters (such as warmup, maximum gradient norm, etc.).